# A Fourier-Jacobi Dirichlet series attached to modular forms of $\textup{SO}(2,4)$

### v1

@@ 8. Representation Numbers of Binary Hermitian Forms
It is shown in [EGM97, Definition 3.4] that this map is well-defined and from [EGM97, Proposition 3.5] induces a map
$\varphi_{f_{i}}:P(f_{i},d)/E(f_{i})\rightarrow\mathcal{R}(d,-\ell),$
@@
where $E(f_{i}):=\left\{\gamma\in\textup{SL}_{2}(\mathcal{O}_{K})\,\,|\,\,\gamma f_{i}\gamma^{*}=f_{i}\right\}$ acting linearly on the set $P(f_{i},d)$. Let now
$\mathcal{R}(f_{i},d):=\left\{h\in\mathcal{O}_{K}/d\mathcal{O}_{K}\mid h\overline{h}\equiv-\ell\pmod{d}\textup{ and }\begin{pmatrix}(h\overline{h}+\ell)/d&h\\ \overline{h}&d\end{pmatrix}\sim f_{i}\right\},$
We then have $\textup{Im}(\varphi_{f_{i}})\subseteq\mathcal{R}(f_{i},d)$ and $r(f_{i},d)=\#\mathcal{R}(f_{i},d)$, where $r(f_{i},d)$ is defined in (31). Our aim is to show that $\varphi_{f_{i}}$ induces a bijection between $P(f_{i},d)/E(f_{i})$ and $\mathcal{R}(f_{i},d)$.
@@
To that end, we form the disjoint union of the domains $P(f_{i},d)/E(f_{i})$ for all $i=1,\cdots,I_{\ell}$, and construct the map
$\varphi_{\ell}:\bigsqcup_{i\in I_{\ell}}P(f_{i},d)/E(f_{i})\longrightarrow\mathcal{R}(d,-\ell)$
@@
such that $\varphi_{\ell}{\mid(P(f_{i},d)/E(f_{i}))}:=\varphi_{f_{i}}$ for all $i=1,\cdots,I_{\ell}$. From [EGM97, Theorem 3.7], the map $\varphi_{\ell}$ is bijective.
In particular, this shows that each $\phi_{f_{i}}$ is injective and surjective onto $\mathcal{R}(f_{i},d)$, hence proving our claim. Hence, if we set $e(f_{i}):=\#E(f_{i})$, we have that $\frac{p(f_{i},d)}{e(f_{i})}=r(f_{i},d)$, and that
$\sum_{i=1}^{I_{\ell}}\frac{p(f_{i},d)}{e(f_{i})}=r(d,-\ell),$
where $r(d,-\ell):=\#\mathcal{R}(d,-\ell)$.
@@ 9. Partial Zeta functions and Latimer’s work
In this Section, we will establish a link between the partial zeta functions $\zeta_{\xi_{i},\mathcal{P}}(s)$ of (14) and partial zeta functions attached to the quaternion algebra $B$. From Proposition 8.2 we have, for any $d\geq 1$, $n(\xi_{i},d)=r(f_{i},d)$, and from the previous discussion $r(f_{i},d)=p(f_{i},d)/e(f_{i})$. Hence,
$\zeta_{\xi_{i},\mathcal{P}}(s)=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}n(\xi_{i},N)N^{-s}=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}r(f_{i},N)N^{-s}=\\ =\frac{1}{e(f_{i})}\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}p(f_{i},N)N^{-s}=\frac{1}{e(f_{i})}\sum_{\begin{subarray}{c}(u,v)\in\mathcal{O}_{K}^{2}\setminus\mathcal{P},\\ <u,v>=\mathcal{O}_{K}\end{subarray}}\frac{1}{f_{i}(u,v)^{s}},$
where the notation $(u,v)\in\mathcal{O}_{K}^{2}\setminus\mathcal{P}$ means that we exclude the pairs $(u,v)$ which give $\gcd(f_{i}(u,v),p)\neq 1$ for $p\in\mathcal{P}$. We now set
$\hat{\zeta}_{\xi_{i},\mathcal{P}}(s):=\zeta_{K,\mathcal{P}}(s)\zeta_{\xi_{i},\mathcal{P}}(s),$ (32)
where $\zeta_{K,\mathcal{P}}(s)$ is the Dedekind zeta function of $K$, with Euler factors above primes in $\mathcal{P}$ removed. We now have the following Lemma.
@@
Assume that $K$ has class number one. Then
$\hat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\frac{1}{e(f_{i})}\sum_{(u,v)\in\mathcal{O}_{K}^{2}\setminus\mathcal{P}}\frac{1}{f_{i}(u,v)^{s}}.$
@@
This follows from factoring the possible common factors of $u$ and $v$ in the second series (compare also with [EGM97, Proposition 4.5]). ∎
To proceed further, we rely on the two papers of Latimer [Lat35, Lat36] on the relation between representation numbers of binary Hermitian forms and ideals in quaternion algebras.
@@
$N(\mathfrak{U}):=\det\begin{pmatrix}g_{11}&g_{12}\\ g_{21}&g_{22}\end{pmatrix}$. (33)
(see [Lat35, Section 4]). In the case when $N(\mathfrak{U})$ is a positive integer, we call $\omega_{1},\omega_{2}$ a proper basis. Following [Lat35, Theorem 1], we will call $\mathfrak{U}$ a regular ideal if it possesses a proper basis. From [Lat35, Lemma 4], if $\mathfrak{U}$ is a regular ideal, every ideal in its (left) class is also regular. Hence, the term regular class is meaningful.
The main theorem in [Lat35, Theorem 3] reads as follows:
@@
There is one-to-one correspondence between regular classes of left ideals in $\mathfrak{o}$ and $\textup{SL}_{2}(\mathcal{O}_{K})$-classes of positive definite binary hermitian forms.
@@
Let us consider a $f_{i}\in\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]\subseteq\Lambda$, and write $f_{i}=\begin{pmatrix}a&b\\ \bar{b}&c\end{pmatrix}$, with $a,c\in\mathbb{Z}$ and $b\in\mathcal{O}_{K}$. By the proof of [Lat35, Theorem 3], a representative of the class corresponding to this form is $\mathcal{L}:=[a,b+\epsilon]$. It suffices to show $\mathcal{L}$ invertible, as then every ideal in its class will be invertible (see [Voi21, Chapter 17]). Recall, $\epsilon^{2}=-k=b\bar{b}-ac<0$. Let $\mathcal{L}^{\prime}:=[a,\bar{b}-\epsilon]$, which is the ideal obtained by taking the standard involution of $B$. We compute
@@
$a\mathcal{O}_{K}+b\mathcal{O}_{K}+\bar{b}\mathcal{O}_{K}+c\mathcal{O}_{K}=\mathcal{O}_{K}.$ (34)
Indeed, if this is true, we will have that $\mathcal{L}\left(\frac{1}{a}\mathcal{L}^{\prime}\right)=\mathfrak{o}$, so it would be enough to show that $\mathfrak{o}=\mathcal{O}_{L}(\mathcal{L})$. Following [Voi21, paragraph 16.5.18], we have
@@
We will make use of the assumption that $f_{i}\in\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]$, and $\ell$ is an integer. Note that from [Shi06a, (2.18)], the first condition is equivalent to
@@
For the other direction, we write $\mathcal{O}_{K}=\mathbb{Z}+\omega\mathbb{Z}=\mathbb{Z}+\overline{\omega}\mathbb{Z}$, with $\omega$ as in (17). We then have
@@
For a (left) ideal $I$ of $\mathfrak{o}$, we set $\mathcal{N}(I):=\#(\mathfrak{o}/I)$. We will now prove the following Lemma, which is given without proof as a footnote in [Lat35, page 441].
@@
If $I$ is a regular left ideal in $\mathfrak{o}$, we have $N(I)^{2}=\mathcal{N}(I)$, where $N(I)$ is the norm of equation (33).
@@
Assume $K$ has class number $1$. For each $i=1,\cdots,h_{1}$, we have
$\hat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\sum_{\begin{subarray}{c}I\in[\mathfrak{U}_{i}]\\ (I,{\mathcal{P}})=1\end{subarray}}N(I)^{-s},$
@@
For this, we rely on [Lat36, Theorem 1]. On the one hand, we have from Lemma 9.1 that
$\hat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\frac{1}{e(f_{i})}\sum_{(u,v)\in\mathcal{O}_{K}^{2}\setminus\mathcal{P}}\frac{1}{f_{i}(u,v)^{s}}=\frac{1}{e(f_{i})}\sum_{\begin{subarray}{c}d=1\\ (d,p)=1\ \forall p\in\mathcal{P}\end{subarray}}^{\infty}q(f_{i},d)d^{-s},$
where $q(f_{i},d):=\#\{(u,v)\in\mathcal{O}_{K}^{2}\backslash\mathcal{P}\mid f_{i}(u,v)=d\}$. From [Lat36, Corollary, p. 443] and because each $f_{i}$ is taken positive definite, we have that for each $d\geq 1$ with $\gcd(d,p)=1$ for all $p\in\mathcal{P}$
@@
We have stated the Proposition above in terms of left ideals in $\mathfrak{o}$, but the presence of the main involution of $B$ allows us to transfer the results also for right-ideals. Given a left ideal $I$ in $\mathfrak{o}$, we can obtain a right ideal $\bar{I}:=\{\bar{a}\,\,|\,\,a\in I\}$, where $\,\bar{}\,$ here denotes the main involution of $B$. Then one obtains a bijection between left and right ideals which preserves norms. Indeed, let us show that $N(\bar{I})=N(I)$. From [Voi21, Section 16.6.14], for any invertible ideal $J$, we have $\bar{J}J=N(J)\mathcal{O}_{R}(J)$, and $J\bar{J}=N(J)\mathcal{O}_{L}(J)$. By setting $J=\bar{I}$ in the first relation and $J=I$ in the second, we obtain
@@
From [Voi21, Lemma 16.6.7], we have $\mathcal{O}_{R}(\bar{I})=\mathcal{O}_{L}(I)$, and hence we obtain that $\textup{nrd}(I)=\textup{nrd}(\bar{I})$. But $\textup{nrd}(I)=N(I)$, from our Lemma 9.4 and [Voi21, Theorem 16.1.3].
@@
Assume $K$ has class number $1$. For each $i=1,\cdots,h_{1}$, we have
$\hat{\zeta}_{\xi_{i},\mathcal{P}}(s)=\sum_{\begin{subarray}{c}I\in[\overline{\mathfrak{U}_{i}}]\\ (I,{\mathcal{P}})=1\end{subarray}}N(I)^{-s},$

### v2

@@ 10. Representation Numbers of Binary Hermitian Forms
It is shown in [Elstrodt, Definition 3.4] that this map is well-defined and from [Elstrodt, Proposition 3.5] induces a map
$\varphi_{f_{i}}:P(f_{i},\mathfrak{m},d)/E(f_{i})\rightarrow\mathcal{R}(\mathfrak{m},d,-\ell),$
@@
where $E(f_{i}):=\left\{\gamma\in\textup{SL}_{2}(\mathcal{O}_{K})\,\,|\,\,\gamma f_{i}\gamma^{*}=f_{i}\right\}$ acting linearly on the set $P(f_{i},\mathfrak{m},d)$. We then have the following Lemma.
Lemma 10.4.
The map $\phi_{f_{i}}$ gives a bijection between $P(f_{i},\overline{\mathfrak{m}},d)/E(f_{i})$ and $\mathcal{R}(f_{i},\mathfrak{m},d)$ of (10.5).
Proof.
@@
First, from the definition above, the image $\phi_{f_{i}}$ is in $\mathcal{R}(f_{i},\mathfrak{m},d)$. We now form the disjoint union of the domains $P(f_{i},\overline{\mathfrak{m}},d)/E(f_{i})$ for all $i=1,\cdots,I_{\ell}$, and construct the map
(10.7) $\varphi_{\ell}:\bigsqcup_{i\in I_{\ell}}P(f_{i},\overline{\mathfrak{m}},d)/E(f_{i})\longrightarrow\mathcal{R}(\overline{\mathfrak{m}},d,-\ell)$
@@
such that $\varphi_{\ell}{\mid(P(f_{i},\overline{\mathfrak{m}},d)/E(f_{i}))}:=\varphi_{f_{i}}$ for all $i=1,\cdots,I_{\ell}$. From [Elstrodt, Theorem 3.7], the map $\varphi_{\ell}$ is bijective.
Since $\phi_{\ell}$ of (10.7) is bijective and $\phi_{f_{i}}=\phi_{\ell}|_{P(f_{i},\mathfrak{m},d)/E(f_{i})}$, we have that $\phi_{f_{i}}$ is injective. We now only need to show that the image is the whole of $\mathcal{R}(f_{i},\mathfrak{m},d)$. Let $h\in\mathcal{R}(f_{i},\mathfrak{m},d)$. Then, since $\phi_{\ell}$ is surjective, there is $(u,v)\in P(f_{j},\overline{\mathfrak{m}},d)$ with $\phi_{\ell}((u,v))=h$ for some $j$. This means that there is $\sigma\in\begin{pmatrix}\mathfrak{m}&\mathfrak{m}\\ \overline{\mathfrak{m}}&\overline{\mathfrak{m}}\end{pmatrix}$ with $\det\sigma=N(\mathfrak{m})$ such that
$\sigma f_{j}\sigma^{*}=\begin{pmatrix}\star&h\\ \overline{h}&dN(\mathfrak{m})\end{pmatrix}$.
But since $h\in R(f_{i},\mathfrak{m},d)$, we also have that $\exists\tau$ with the same properties such that
$\tau f_{j}\tau^{*}=\begin{pmatrix}\star&h\\ \overline{h}&dN(\mathfrak{m})\end{pmatrix}$.
The $\star$ entries are the same because the matrix has determinant $\ell N(\mathfrak{m})^{2}$. Hence $\tau^{-1}\sigma f_{j}(\tau^{-1}\sigma)^{*}=f_{i}$. But we can now verify $\tau^{-1}\sigma\in\textup{SL}_{2}(\mathcal{O}_{K})$, which forces $i=j$. Hence, the Lemma follows. ∎
@@ 11. Partial Zeta functions and Latimer’s work
In this Section, we will establish a link between the partial zeta functions $\zeta_{\mathcal{Q}}(\xi_{\mathfrak{m}i};s)$ of (6.4) and partial zeta functions attached to the quaternion algebra $B$. From Proposition 10.3 we have, for any $d\geq 1$, $n(\xi_{\mathfrak{m},i},d)=r(f_{i},\mathfrak{m},d)$, and from the previous discussion $r(f_{i},\mathfrak{m},d)=p(f_{i},\mathfrak{m},d)/e(f_{i})$. Hence,
$\zeta_{\mathcal{Q}}(\xi_{\mathfrak{m},i};s)=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{Q}\end{subarray}}^{\infty}n(\xi_{\mathfrak{m},i};N)N^{-s}=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{Q}\end{subarray}}^{\infty}r(f_{i},\mathfrak{m},N)N^{-s}=\\ =\frac{1}{e(f_{i})}\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{Q}\end{subarray}}^{\infty}p(f_{i},\mathfrak{m},N)N^{-s}=\frac{1}{e(f_{i})}N(\mathfrak{m})^{s}\sum_{\begin{subarray}{c}(u,v)\in\mathfrak{m}^{2}\setminus\mathcal{Q},\\ <u,v>=\mathfrak{m}\end{subarray}}\frac{1}{f_{i}(u,v)^{s}},$
where the notation $(u,v)\in\mathfrak{m}^{2}\setminus\mathcal{Q}$ means that we exclude the pairs $(u,v)$ which give $\gcd(f_{i}(u,v),p)\neq 1$ for $p\in\mathcal{Q}$. We now set
(11.1) $\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s):=\sum_{\mathfrak{m}\in\textup{Cl}(K)}\zeta_{\mathfrak{m},\mathcal{Q}}(s)\zeta_{\mathcal{Q}}(\xi_{\mathfrak{m},i};s).$
In particular, equation (8.4) now reads
(11.2) $D(\bm{F},\bm{P};s)=\sum_{i=1}^{h_{1}}\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s-k+3)D_{\mathcal{Q}}(F,\iota(\xi_{i});s).$
We now have the following Lemma.
@@
With notation as above we have,
$\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s)=\frac{1}{e(f_{i})}\sum_{(u,v)\in\mathcal{O}_{K}^{2}\setminus\mathcal{Q}}\frac{1}{f_{i}(u,v)^{s}}.$
@@
This is [Elstrodt, Proposition 4.5]. ∎
To proceed further, we rely on the two papers of Latimer [Latimer1, Latimer2] on the relation between representation numbers of binary Hermitian forms and ideals in quaternion algebras.
@@
(11.3) $N(\mathfrak{U}):=\det\begin{pmatrix}g_{11}&g_{12}\\ g_{21}&g_{22}\end{pmatrix}$.
(see [Latimer1, Section 4]). In the case when $N(\mathfrak{U})$ is a positive integer, we call $\omega_{1},\omega_{2}$ a proper basis. Following [Latimer1, Theorem 1], we will call $\mathfrak{U}$ a regular ideal if it possesses a proper basis. From [Latimer1, Lemma 4], if $\mathfrak{U}$ is a regular ideal, every ideal in its (left) class is also regular. Hence, the term regular class is meaningful.
The main theorem in [Latimer1, Theorem 3] reads as follows:
@@
There is one-to-one correspondence between regular classes of left ideals in $\mathfrak{o}$ and $\textup{SL}_{2}(\mathcal{O}_{K})$-classes of positive definite binary hermitian forms of discriminant $\ell$.
@@
Let us consider a $f_{i}\in\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]\subseteq\Lambda$, and write $f_{i}=\begin{pmatrix}a&b\\ \bar{b}&c\end{pmatrix}$, with $a,c\in\mathbb{Z}$ and $b\in\mathcal{O}_{K}$. By the proof of [Latimer1, Theorem 3], a representative of the class corresponding to this form is $\mathcal{L}:=[a,b+\epsilon]$. It suffices to show $\mathcal{L}$ invertible, as then every ideal in its class will be invertible (see [voight, Chapter 17]). Recall, $\epsilon^{2}=-k=b\bar{b}-ac<0$.
Let $\mathcal{L}^{\prime}:=[a,\bar{b}-\epsilon]$, which is the ideal obtained by taking the standard involution of $B$. We compute
@@
(11.4) $a\mathcal{O}_{K}+b\mathcal{O}_{K}+\bar{b}\mathcal{O}_{K}+c\mathcal{O}_{K}=\mathcal{O}_{K}.$
Indeed, if this is true, we will have that $\mathcal{L}\left(\frac{1}{a}\mathcal{L}^{\prime}\right)=\mathfrak{o}$, so it would be enough to show that $\mathfrak{o}=\mathcal{O}_{L}(\mathcal{L})$. Following [voight, paragraph 16.5.18], we have
@@
We will make use of the assumption that $f_{i}\in\Lambda^{+}\left[\ell,2^{-1}\mathbb{Z}\right]$, and $\ell$ is an integer. Note that from [shimura_diophantine_2006, (2.18)], the first condition is equivalent to
@@
For the other direction, we write $\mathcal{O}_{K}=\mathbb{Z}+\omega\mathbb{Z}=\mathbb{Z}+\overline{\omega}\mathbb{Z}$, with $\omega$ as in (4.1). We then have
@@
For a (left) ideal $I$ of $\mathfrak{o}$, we set $\mathcal{N}(I):=\#(\mathfrak{o}/I)$. We will now prove the following Lemma, which is given without proof as a footnote in [Latimer1, page 441].
@@
If $I$ is a regular left ideal in $\mathfrak{o}$, we have $N(I)^{2}=\mathcal{N}(I)$, where $N(I)$ is the norm of equation (11.3).
@@
For each $i=1,\cdots,h_{1}$, we have
$\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s)=\sum_{\begin{subarray}{c}I\in[\mathfrak{U}_{i}]\\ (I,{\mathcal{Q}})=1\end{subarray}}N(I)^{-s},$
@@
For this, we rely on [Latimer2, Theorem 1]. On the one hand, we have from Lemma 11.1 that
$\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s)=\frac{1}{e(f_{i})}\sum_{(u,v)\in\mathcal{O}_{K}^{2}\setminus\mathcal{Q}}\frac{1}{f_{i}(u,v)^{s}}=\frac{1}{e(f_{i})}\sum_{\begin{subarray}{c}d=1\\ (d,p)=1\ \forall p\in\mathcal{Q}\end{subarray}}^{\infty}q(f_{i},d)d^{-s},$
where $q(f_{i},d):=\#\{(u,v)\in\mathcal{O}_{K}^{2}\backslash\mathcal{Q}\mid f_{i}(u,v)=d\}$. From [Latimer2, Corollary, p. 443] and because each $f_{i}$ is taken positive definite, we have that for each $d\geq 1$ with $\gcd(d,p)=1$ for all $p\in\mathcal{Q}$
@@
We have stated the Proposition above in terms of left ideals in $\mathfrak{o}$, but the presence of the main involution of $B$ allows us to transfer the results also for right-ideals. Given a left ideal $I$ in $\mathfrak{o}$, we can obtain a right ideal $\bar{I}:=\{\bar{a}\,\,|\,\,a\in I\}$, where $\,\bar{}\,$ here denotes the main involution of $B$. Then one obtains a bijection between left and right ideals which preserves norms. Indeed, let us show that $N(\bar{I})=N(I)$. From [voight, Section 16.6.14], for any invertible ideal $J$, we have $\bar{J}J=N(J)\mathcal{O}_{R}(J)$, and $J\bar{J}=N(J)\mathcal{O}_{L}(J)$. By setting $J=\bar{I}$ in the first relation and $J=I$ in the second, we obtain
@@
From [voight, Lemma 16.6.7], we have $\mathcal{O}_{R}(\bar{I})=\mathcal{O}_{L}(I)$, and hence we obtain that $\textup{nrd}(I)=\textup{nrd}(\bar{I})$. But $\textup{nrd}(I)=N(I)$, from our Lemma 11.4 and [voight, Theorem 16.1.3].
@@
For each $i=1,\cdots,h_{1}$, we have
$\widehat{\zeta}_{\mathcal{Q}}(\xi_{i};s)=\sum_{\begin{subarray}{c}I\in[\overline{\mathfrak{U}_{i}}]\\ (I,{\mathcal{Q}})=1\end{subarray}}N(I)^{-s},$